If the mass is tripled to 15kg*3=45kg, what is the new change of motion
1
Expert's answer
2020-12-21T10:40:49-0500

I assume that the particle with the mass "m_1 = 15kg" is moving under the constant force "F". Then, according to the second Newton's law, the acceleration of the particle is:


"a_1 = \dfrac{F}{m_1}"

If we triple the mass "m_2 = 3m_1 = 45kg" and remain force the same, then the new acceleration will be:


"a_2 = \dfrac{F}{m_2} = \dfrac{F}{3m_1} = \dfrac{a_1}{3}"

Answer. The change of motion is the following: acceleration decreases by 3.